# Thermodynamic and observational implications of black holes in toroidal geometry

### v1

@@ Abstract
We investigate the thermodynamic and observational implications for the charged torus-like black holes, a class of solutions distinct from the classical Schwarzschild black holes. We explicitly derive the fundamental thermodynamic properties, such as heat capacity, P-V diagram, isothermal compressibility, Helmholtz free energy, and Gibbs free energy, under different entropy models. We find that only the exponential corrected entropy demonstrates multiple phase transitions, which we validate with the Ricci Scalar divergence obtained from the Ruppeiner formalism. This indicates that exponential corrected entropy is more sensitive to BH’s microstructure as compared to the Hawking-Bekenstein and Rènyi entropy models. In addition, we study the sparsity and emission rates of Hawking radiation, demonstrating that exponential correction entropy yields more consistent and stable behavior. In our observational analysis, we graphically demonstrate the behavior of redshift, blueshift, and gravitational shift, and identify specific conditions where the photon sphere radius exceeds the innermost stable circular orbit radius, which depends on the values of parameters such as electric charge and cosmological constant. The novel insight of this work is that despite this violation, our computed redshift, blueshift, and gravitational shifts fall within the range of the observational data of NGC 4258 and UGC 3789.
@@ I Introduction
Black holes (BHs) have become one of the most important topics in the theory of general relativity, which provides a very unique framework where the concepts of thermodynamics, gravity, and observational astrophysics align. Hawking and Bekenstein’s foundational work on BH thermodynamics has established a significant link between gravitational theory and quantum mechanics Hawking:1974rv ; Bekenstein:1973ur ; Bekenstein:1974ax ; Hawking:1975vcx ; Bekenstein:1972tm ; Bardeen:1973gs . In this framework, BHs follows the laws of thermodynamics, where surface gravity corresponds to temperature (Hawking temperature), the area of the event horizon signifies entropy and exhibiting phase transition akin to those in conventional thermodynamic system Bardeen:1973gs . Analytical frameworks like PVP-V criticality Chamblin:1999tk ; Chamblin:1999hg ; Kubiznak:2012wp ; Kubiznak:2016qmn , stability analysis utilizing heat capacity, Gibbs free energy (GFE) and Helmholtz free energy (HFE) Rodrigues:2022qdp ; Simovic:2023yuv ; Ali:2019myr , provide significant insights on BH characteristics and their thermodynamic behavior. For example, a BH is considered locally and globally stable if the heat capacity, HFE, and GFE demonstrate positive behavior and vice versa (for further details regarding local and global stability of BHs, check Refs. Rodrigues:2022zph ; Dehyadegari:2017flm ; Wei:2020poh ; Liu:2014gvf ; Dehyadegari:2020tau ).
@@
In this context, thermodynamic geometry provides an effective framework to study the BH phase transition by introducing various thermodynamic metrics, and these metrics are constructed in terms of entropy, which gives us valuable insight regarding the phase transition when the divergence of Ricci curvature scalar determined from this metric formalism aligns with the zero point or divergence of heat capacity. The metric formalism introduced by the Weinhold Weinhold:1975xej , based on equilibrium state space, which was later extended by the Ruppiener Ruppeiner:2008kd ; Ruppeiner:1995zz , whose formalism is closely linked with the Weinhold’s formalism (for other thermodynamic geometry formalism and their constructions check Refs. NaveenaKumara:2020biu ; Quevedo:2008ry ; Akbar:2011qw ; Hendi:2015cka ; Sarkar:2006tg ; Bhattacharya:2019qxe ; Hendi:2015xya ; Banerjee:2016nse ), it reveals valuable insights regarding the microscopic interaction and phase transitions. Although research on non-spherical BH configurations has increased, toroidal or torus-like BH have not been investigated as extensively as spherically symmetric ones. The main reason is that they have very intricate topology and horizon geometry, which makes them difficult to analyze. These geometries naturally arise within AdS spacetime and offer a framework for studying the thermodynamic and observational characteristics. In torus-like spacetimes, entropy correction plays a significant role because the traditional entropy (Hawking-Bekenstein entropy) formulated under classical assumptions may not fully incorporate quantum or topological complexities. To better describe these systems, Rènyi Odintsov:2023qfj and exponential corrected entropy ghosht introduce generic statistical or quantum modifications that may be fundamentally significant.
@@
Furthermore, at a constant temperature, Hawking radiation produces sparse and discrete quanta which deviate from the continuous emission pattern of classical blackbody radiation Page:2004xp ; Gray:2015pma . The nature of Hawking radiation significantly impacts both its detectability and quantum information aspects. In recent years, strong evidence supporting the existence of BHs has emerged from the recent observational phenomena, including the stars’ orbital motion near the Galactic center Ghez:2008ms ; Morris:2012uz ; Eckart:1996zz ; Gillessen:2008qv , imaging of supermassive BHs in M87 and the Milky Way EventHorizonTelescope:2019ths ; EventHorizonTelescope:2022xqj , and the gravitational wave observation by LIGO-Virgo LIGOScientific:2016aoc ; LIGOScientific:2016sjg . These findings have encouraged the formulation of methods to determine the BH parameter by using observational data, for example, to determine properties like mass and spin, one widely used approach is to study the frequency shifts (red shift and blueshift) of photons emitted by orbiting particles, originally applied in Kerr geometries Herrera-Aguilar:2015kea ; Banerjee:2022him ; Momennia:2023lau . Subsequent development has extended this methodology to a wider class of BH solutions, including Kerr-Newmann dS Kraniotis:2019ked , higher dimensional Sharif:2016rhs , regular Becerril:2016qxf , polymerized types Fu:2022afk , and has also been employed to compute the Hubble constant Momennia:2023lau . Numerous works have been done to apply general relativity to real astrophysical systems, for example, active galactic nuclei (AGN) with water maser disks, for calculating the ratio between the BH mass and distance Nucamendi:2020tov ; Villalobos-Ramirez:2022xic ; Villaraos:2022euo ; Gonzalez-Juarez:2022pya . Recently, the concept of redshift rapidity, have made it possible to distinguish distance from mass in Reisner-nordstörm (RN) and Schwarzschild spacetimes with the help of observational data Momennia:2023svi ; Morales-Herrera:2024zbr . The values of parameters such as electric charge and spin still exhibit ambiguity, specifically in different models of the modified theory of gravity Ghez:2008ms ; Morris:2012uz ; Eckart:1996zz ; Gillessen:2008qv ; EventHorizonTelescope:2019ths ; EventHorizonTelescope:2022xqj ; LIGOScientific:2016aoc ; LIGOScientific:2016sjg . When relativistic formalisms are extended to a general spherically symmetric spacetime, it allows us to investigate the impacts of additional parameters predicted by modified theories of gravity and analyze their deviation from Einstein’s theory Herrera-Aguilar:2015kea ; Banerjee:2022him ; Momennia:2023lau . We extended this general spherical symmetric formalism to discuss the frequency shifts (redshift, blueshift, and gravitational shift) for the torus-like BH solution, which has toroidal geometry. In addition, studying frequency shifts, which is commonly done for spherical BHs, is equally crucial for torus-like geometries, as it helps in examining strong-field lensing, accretion disk behavior, and deviations in trajectories of photons and orbital stability. The resulting deviations may lead to observable astrophysical signals, establishing frequency shifts, as an effective method for identifying non-traditional horizon geometries. Moreover, we use the cosmological constant instead of using a scaling parameter to analyze its impact on the frequency shifts.
@@ III Thermodynamics Quantities of Torus-like BH Through the Corrected Entropies
where $\mathcal{A}=4\pi r^{2}_{\mathrm{e}}$ presents the area of the horizon and $G$ is the gravitational constant. The standard semiclassical expression for BH entropy is believed to be altered by quantum gravity, notably through exponential corrections that arise naturally in frameworks like non-local gravity, string theory, and quantum statistical approaches. Such exponential terms capture non-perturbative contributions and reveal crucial details about the microscopic framework of the spacetime. Furthermore, these terms may enhance the study of thermodynamic stability, phase transitions, and singularity resolution, effectively associating classical BH thermodynamics with foundational features of quantum gravity. In this perspective, by following the same methodology as described in Ref. ghosht , the expression for the exponential corrected entropy can be written as
@@
where $l^{2}_{p}=~1,~\mathcal{A}=4\pi{r^{2}_{\mathrm{e}}}$ and $\nabla=\frac{\ln{2}}{2\pi}$. By inserting these values, one can easily derive the following form of exponential corrected entropy
$S_{\mathrm{Exp}}=\exp\left(-{\pi}r^{2}_{\mathrm{e}}\right)+{\pi}r^{2}_{\mathrm{e}}\,.$ (9)
@@
where $\alpha$ depicts the non-extensive parameter. The range $0<\alpha<1$, ensures that the Rényi entropy $S_{\mathrm{R}}$ is mathematically consistent and thermodynamically viable, whereas values beyond this limit lead to an undefined form because of the entropy’s convex behavior (for more details in terms of considering non-extensive parameter $\alpha=0$ and other non-extensive entropy framewoks, see Refs. Odintsov:2023qfj ; Nojiri:2022aof ; Nojiri:2022dkr ; Nojiri:2022sfd ; Elizalde:2025iku ; Odintsov:2023vpj ; Nojiri:2024zdu ).
@@ III.1 Thermodynamic Mass
$M(S_{\mathrm{HB}},~P,~Q)=\frac{2\sqrt{\pi}\left(2PS^{2}+3Q^{2}\right)}{3\sqrt{S}}\,.$ (11)
@@
$M(S_{\mathrm{Exp}},~P,~Q)$ $=$ $\frac{2^{3/4}\sqrt{\pi}\left\{4P(S-1)+3Q^{2}\right\}}{3\sqrt[4]{S-1}}\,.$ (12)
@@
$M(S_{\mathrm{R}},~P,~Q)=\frac{2\sqrt{\pi}\sqrt{\alpha}\left\{\frac{2P\left(e^{\alpha S}-1\right)^{2}}{\alpha^{2}}+3Q^{2}\right\}}{3\sqrt{e^{\alpha S}-1}}\,.$ (13)
@@ III.2 Temperature
$T_{\mathrm{HB}}=\frac{\partial M_{\mathrm{HB}}}{\partial S_{\mathrm{HB}}}$ $=$ $\frac{2\sqrt{\pi}\left(2PS^{2}+3Q^{2}\right)}{3\sqrt{S}}\,$,
$V_{\mathrm{HB}}=\frac{\partial M_{\mathrm{HB}}}{\partial P}$ $=$ $\frac{4}{3}\sqrt{\pi}S^{3/2}\,$,
$\phi_{\mathrm{HB}}=\frac{\partial M_{\mathrm{HB}}}{\partial Q}$ $=$ $\frac{4\sqrt{\pi}Q}{\sqrt{S}}\,.$ (14)
@@
$T_{\mathrm{Exp}}=\frac{\partial M_{\mathrm{Exp}}}{\partial S_{\mathrm{Exp}}}$ $=$ $\frac{2^{3/4}\sqrt{\pi}\left\{4P(S-1)+3Q^{2}\right\}}{3\sqrt[4]{S-1}}\,$,
$V_{\mathrm{Exp}}=\frac{\partial M_{\mathrm{Exp}}}{\partial P}$ $=$ $\frac{4}{3}2^{3/4}\sqrt{\pi}(S-1)^{3/4}\,$,
$\phi_{\mathrm{Exp}}=\frac{\partial M_{\mathrm{Exp}}}{\partial Q}$ $=$ $\frac{2\ 2^{3/4}\sqrt{\pi}Q}{\sqrt[4]{S-1}}\,.$ (16)
@@
$T_{\mathrm{R}}=\frac{\partial M_{\mathrm{R}}}{\partial S_{\mathrm{R}}}$ $=$ $\frac{\sqrt{\pi}e^{\alpha S}\left\{2P\left(e^{\alpha S}-1\right)^{2}-\alpha^{2}Q^{2}\right\}}{\sqrt{\alpha}\left(e^{\alpha S}-1\right)^{3/2}}\,$,
$V_{\mathrm{R}}=\frac{\partial M_{\mathrm{R}}}{\partial V_{\mathrm{R}}}$ $=$ $\frac{4\sqrt{\pi}\left(e^{\alpha S}-1\right)^{3/2}}{3\alpha^{3/2}}\,$,
$\phi_{\mathrm{R}}=\frac{\partial M_{R}}{\partial Q}$ $=$ $\frac{4\sqrt{\pi}\sqrt{\alpha}Q}{\sqrt{e^{\alpha S}-1}}\,.$ (18)

### v2

@@ Abstract
We investigate the thermodynamic and observational implications for the charged torus-like black holes, a class of solutions distinct from the classical Schwarzschild black holes. We explicitly derive the fundamental thermodynamic properties, such as heat capacity, P-V diagram, isothermal compressibility, Helmholtz free energy, and Gibbs free energy, under different entropy models. We find that only the exponential corrected entropy demonstrates multiple phase transitions, which we validate with the Ricci Scalar divergence obtained from the Ruppeiner formalism. This indicates that exponential corrected entropy is more sensitive to BH’s microstructure as compared to the Hawking-Bekenstein and Rènyi entropy models within the non-spherical (toroidal) horizon. In addition, we study the sparsity and emission rates of Hawking radiation, demonstrating that exponential correction entropy yields more consistent and stable behavior. In our observational analysis, we graphically demonstrate the behavior of redshift, blueshift, and gravitational shift, and identify specific conditions where the photon sphere radius exceeds the innermost stable circular orbit radius, which depends on the values of parameters such as electric charge and cosmological constant. The novel insight of this work is that despite this violation, our computed redshift, blueshift, and gravitational shifts fall within the range of the observational data of NGC 4258 and UGC 3789.
@@ I Introduction
Black holes (BHs) have become one of the most important topics in the theory of general relativity, which provides a very unique framework where the concepts of thermodynamics, gravity, and observational astrophysics align. Hawking and Bekenstein’s foundational work on BH thermodynamics has established a significant link between gravitational theory and quantum mechanics Hawking:1974rv ; Bekenstein:1973ur ; Bekenstein:1974ax ; Hawking:1975vcx ; Bekenstein:1972tm ; Bardeen:1973gs . In this framework, BHs follows the laws of thermodynamics, where surface gravity corresponds to temperature (Hawking temperature), the area of the event horizon signifies entropy and exhibiting phase transition akin to those in conventional thermodynamic system Bardeen:1973gs . Analytical frameworks like PVP-V criticality Chamblin:1999tk ; Chamblin:1999hg ; Kubiznak:2012wp ; Kubiznak:2016qmn ; Karch:2015rpa ; Mancilla:2024spp , stability analysis utilizing heat capacity, Gibbs free energy (GFE) and Helmholtz free energy (HFE) Rodrigues:2022qdp ; Simovic:2023yuv ; Ali:2019myr , provide significant insights on BH characteristics and their thermodynamic behavior. For example, a BH is considered locally and globally stable if the heat capacity, HFE, and GFE demonstrate positive behavior and vice versa (for further details regarding local and global stability of BHs, check Refs. Rodrigues:2022zph ; Dehyadegari:2017flm ; Wei:2020poh ; Liu:2014gvf ; Dehyadegari:2020tau ).
@@
In this context, thermodynamic geometry provides an effective framework to study the BH phase transition by introducing various thermodynamic metrics, and these metrics are constructed in terms of entropy, which gives us valuable insight regarding the phase transition when the divergence of Ricci curvature scalar determined from this metric formalism aligns with the zero point (ZP) or divergence of heat capacity. The metric formalism introduced by the Weinhold Weinhold:1975xej , based on equilibrium state space, which was later extended by the Ruppiener Ruppeiner:2008kd ; Ruppeiner:1995zz , whose formalism is closely linked with the Weinhold’s formalism (for other thermodynamic geometry formalism and their constructions check Refs. NaveenaKumara:2020biu ; Quevedo:2008ry ; Akbar:2011qw ; Hendi:2015cka ; Sarkar:2006tg ; Bhattacharya:2019qxe ; Hendi:2015xya ; Banerjee:2016nse ), it reveals valuable insights regarding the microscopic interaction and phase transitions. Although research on non-spherical BH configurations has increased, toroidal or torus-like BHs have not been investigated as extensively as spherically symmetric ones. The main reason is that they have very intricate topology and horizon geometry, which makes them difficult to analyze. These geometries naturally arise within AdS spacetime and offer a framework for studying the thermodynamic and observational characteristics. Departing from previous analyses, which focus on the hyperbolic and planar horizons, our toroidal framework features a novel dual periodic topology, providing modified thermodynamic curvature characteristics. In torus-like spacetimes, entropy correction plays a significant role because the traditional entropy (Hawking-Bekenstein entropy) formulated under classical assumptions may not fully incorporate quantum or topological complexities. To better describe these systems, Rènyi Odintsov:2023qfj and exponential corrected entropy ghosht introduce generic statistical or quantum modifications that may be fundamentally significant.
@@
Furthermore, at a constant temperature, Hawking radiation produces sparse and discrete quanta which deviate from the continuous emission pattern of classical blackbody radiation Page:2004xp ; Gray:2015pma . The nature of Hawking radiation significantly impacts both its detectability and quantum information aspects. Moreover, we also address the Hawking radiation by exploring emission rates and sparsity, which elucidate the effects of entropy corrections on BH evaporation. In recent years, strong evidence supporting the existence of BHs has emerged from the recent observational phenomena, including the stars’ orbital motion near the Galactic center Ghez:2008ms ; Morris:2012uz ; Eckart:1996zz ; Gillessen:2008qv , imaging of supermassive BHs in M87 and the Milky Way EventHorizonTelescope:2019ths ; EventHorizonTelescope:2022xqj , and the gravitational wave observation by LIGO-Virgo LIGOScientific:2016aoc ; LIGOScientific:2016sjg . These findings have encouraged the formulation of methods to determine the BH parameter by using observational data, for example, to determine properties like mass and spin, one widely used approach is to study the frequency shifts (red shift and blueshift) of photons emitted by orbiting particles, originally applied in Kerr geometries Herrera-Aguilar:2015kea ; Banerjee:2022him ; Momennia:2023lau . Subsequent development has extended this methodology to a wider class of BH solutions, including Kerr-Newmann dS Kraniotis:2019ked , higher dimensional Sharif:2016rhs , regular Becerril:2016qxf , polymerized types Fu:2022afk , and has also been employed to compute the Hubble constant Momennia:2023lau . Numerous works have been done to apply general relativity to real astrophysical systems, for example, active galactic nuclei (AGN) with water maser disks, for calculating the ratio between the BH mass and distance Nucamendi:2020tov ; Villalobos-Ramirez:2022xic ; Villaraos:2022euo ; Gonzalez-Juarez:2022pya . Recently, the concept of redshift rapidity, have made it possible to distinguish distance from mass in Reisner-nordstörm (RN) and Schwarzschild spacetimes with the help of observational data Momennia:2023svi ; Morales-Herrera:2024zbr . The values of parameters such as electric charge and spin still exhibit ambiguity, specifically in different models of the modified theory of gravity Ghez:2008ms ; Morris:2012uz ; Eckart:1996zz ; Gillessen:2008qv ; EventHorizonTelescope:2019ths ; EventHorizonTelescope:2022xqj ; LIGOScientific:2016aoc ; LIGOScientific:2016sjg . When relativistic formalisms are extended to a general spherically symmetric spacetime, it allows us to investigate the impacts of additional parameters predicted by modified theories of gravity and analyze their deviation from Einstein’s theory Herrera-Aguilar:2015kea ; Banerjee:2022him ; Momennia:2023lau . We extended this general spherical symmetric formalism to discuss the frequency shifts (redshift, blueshift, and gravitational shift) for the torus-like BH solution, which has toroidal geometry. In addition, studying frequency shifts, which is commonly done for spherical BHs, is equally crucial for torus-like geometries, as it helps in examining strong-field lensing, accretion disk behavior, and deviations in trajectories of photons and orbital stability. The resulting deviations may lead to observable astrophysical signals, establishing frequency shifts, as an effective method for identifying non-traditional horizon geometries. We also demonstrate the observational significance of our framework for these types of BHs by comparing its predicted frequency shifts with measurements from megamaser systems.
@@ III Thermodynamics Quantities of Torus-like BH Through the Corrected Entropies
where $\mathcal{A}=4\pi^{2}r^{2}_{\mathrm{e}}$ presents the area of the horizon Feng:2021vey , which is different from the spherical case, and $G$ is the gravitational constant. We mention here that this distinction is quite crucial for maintaining consistent thermodynamic parameters, such as entropy and volume, since the geometric factor $4\pi^{2}$ corresponds to the $S^{\star}\times S^{\star}$ horizon topology Ali:2023pyv that we employ consistently in our thermodynamic analysis. The standard semiclassical expression for BH entropy is believed to be altered by quantum gravity, notably through exponential corrections that arise naturally in frameworks like non-local gravity, string theory, and quantum statistical approaches. Such exponential terms capture non-perturbative contributions and reveal crucial details about the microscopic framework of the spacetime. Furthermore, these terms may enhance the study of thermodynamic stability, phase transitions, and singularity resolution, effectively associating classical BH thermodynamics with foundational features of quantum gravity. In this perspective, by following the same methodology as described in Ref. ghosht , the expression for the exponential corrected entropy can be written as
@@
where $l^{2}_{p}=~1,~\mathcal{A}=4\pi^{2}{r^{2}_{\mathrm{e}}}$ and $\nabla=\frac{\ln{2}}{2\pi}$. By inserting these values, one can easily derive the following form of exponential corrected entropy
$S_{\mathrm{Exp}}=\exp\left(-\pi^{2}r^{2}_{\mathrm{e}}\right)+\pi^{2}r^{2}_{\mathrm{e}}\,.$ (9)
@@
where $\alpha$ depicts the non-extensive parameter. The range $0<\alpha<1$, ensures that the Rényi entropy $S_{\mathrm{R}}$ is mathematically consistent and thermodynamically viable, whereas values beyond this limit lead to an undefined form because of the entropy’s convex behavior (for more details in terms of considering non-extensive parameter $\alpha=0$ and other non-extensive entropy framewoks, see Refs. Odintsov:2023qfj ; Nojiri:2022aof ; Nojiri:2022dkr ; Nojiri:2022sfd ; Elizalde:2025iku ; Odintsov:2023vpj ; Nojiri:2024zdu ). Thereby, by comparing these above mention entropy models enables us to examine how quantum correction and non-extensive effects modifies the thermodynamic phase structure of BHs with complex topology.
@@ III.1 Thermodynamic Mass
$M(S_{\mathrm{HB}},~P,~Q)=\frac{2\pi}{3\sqrt{S}}\left(\frac{2PS^{2}}{\pi^{2}}+3Q^{2}\right)\,.$ (11)
@@
$M(S_{\mathrm{Exp}},~P,~Q)$ $=$ $\frac{2^{3/4}\pi}{3\sqrt[4]{S-1}}\left(\frac{4P(S-1)}{\pi^{2}}+3Q^{2}\right)\,.$ (12)
@@
$M(S_{\mathrm{R}},~P,~Q)=\frac{2\pi\sqrt{\alpha}}{3\sqrt{e^{\alpha S}-1}}\left\{\frac{2P\left(e^{\alpha S}-1\right)^{2}}{\pi^{2}\alpha^{2}}+3Q^{2}\right\}\,.$ (13)
@@ III.2 Temperature
$T_{\mathrm{HB}}=\frac{\partial M_{\mathrm{HB}}}{\partial S_{\mathrm{HB}}}$ $=$ $\frac{2P\sqrt{S}}{\pi}-\frac{\pi Q^{2}}{S^{3/2}}\,$,
$V_{\mathrm{HB}}=\frac{\partial M_{\mathrm{HB}}}{\partial P}$ $=$ $\frac{4S^{3/2}}{3\pi}\,$,
$\phi_{\mathrm{HB}}=\frac{\partial M_{\mathrm{HB}}}{\partial Q}$ $=$ $\frac{4\pi Q}{\sqrt{S}}\,.$ (14)
@@
$T_{\mathrm{Exp}}=\frac{\partial M_{\mathrm{Exp}}}{\partial S_{\mathrm{Exp}}}$ $=$ $\frac{4P(S-1)-\pi^{2}Q^{2}}{2\sqrt[4]{2}\pi(S-1)^{5/4}}\,$,
$V_{\mathrm{Exp}}=\frac{\partial M_{\mathrm{Exp}}}{\partial P}$ $=$ $\frac{4\ \times 2^{3/4}(S-1)^{3/4}}{3\pi}\,$,
$\phi_{\mathrm{Exp}}=\frac{\partial M_{\mathrm{Exp}}}{\partial Q}$ $=$ $\frac{2\ 2^{3/4}\pi Q}{\sqrt[4]{S-1}}\,.$ (16)
@@
$T_{\mathrm{R}}=\frac{\partial M_{\mathrm{R}}}{\partial S_{\mathrm{R}}}$ $=$ $\frac{e^{\alpha S}\left(2P\left(e^{\alpha S}-1\right)^{2}-\pi^{2}\alpha^{2}Q^{2}\right)}{\pi\sqrt{\alpha}\left(e^{\alpha S}-1\right)^{3/2}}\,$,
$V_{\mathrm{R}}=\frac{\partial M_{\mathrm{R}}}{\partial V_{\mathrm{R}}}$ $=$ $\frac{4\left(e^{\alpha S}-1\right)^{3/2}}{3\pi\alpha^{3/2}}\,$,
$\phi_{\mathrm{R}}=\frac{\partial M_{R}}{\partial Q}$ $=$ $\frac{4\pi\sqrt{\alpha}Q}{\sqrt{e^{\alpha S}-1}}\,.$ (18)